OpenStudy (anonymous):

what is the solution of |-1+x|=5?

OpenStudy (tkhunny):

Well, it must be either of these: x - 1 = 5 or x - 1 = -5 You tell me which, or both, and why!

OpenStudy (anonymous):

would it be x-1=-5?

OpenStudy (tkhunny):

Yes, and the other one. Solve them both and see if they work in the ORIGINAL equation.

OpenStudy (anonymous):

Absolute value is a measure of distance. Because distance can be measured in any direction, it's both.

OpenStudy (anonymous):

Thank yall!